OpenStudy (anonymous):

What is the value of the expression? 12.4 + (–2 + 1.6) A. 11 B. 12 C. 15 D. 16

OpenStudy (jango_in_dtown):

12.4 + (–2 + 1.6)=12.4+1.4=13.4=52

OpenStudy (triciaal):